Problem 15-03

Nash Company has two classes of capital stock outstanding: 8%, $20 par preferred and $5 par common. At December 31, 2020, the following accounts were included in stockholders’ equity.

Preferred Stock, 148,200 shares $ 2,964,000
Common Stock, 1,991,000 shares 9,955,000
Paid-in Capital in Excess of Par—Preferred Stock 207,000
Paid-in Capital in Excess of Par—Common Stock 26,478,000
Retained Earnings 4,495,000


The following transactions affected stockholders’ equity during 2021.

Jan. 1 29,200 shares of preferred stock issued at $23 per share.
Feb. 1 46,100 shares of common stock issued at $21 per share.
June 1 2-for-1 stock split (par value reduced to $2.50).
July 1 28,900 shares of common treasury stock purchased at $10 per share. Nash uses the cost method.
Sept. 15 9,200 shares of treasury stock reissued at $12 per share.
Dec. 31 The preferred dividend is declared, and a common dividend of 53¢ per share is declared.
Dec. 31 Net income is $2,138,000.

Prepare the stockholders’ equity section for Nash Company at December 31, 2021. (Enter account name only and do not provide descriptive information.)

NASH COMPANY
Stockholders’ Equity

Entry field with incorrect answer now contains modified data December 31, 2021For the Year Ended December 31, 2021For the Quarter Ended December 31, 2021

Entry field with correct answer Additional Paid-in CapitalCapital StockCurrent AssetsCurrent LiabilitiesIntangible AssetsLong-term InvestmentsLong-term LiabilitiesProperty, Plant and EquipmentStockholders' EquityTotal AssetsTotal Capital StockTotal Current AssetsTotal Current LiabilitiesTotal Intangible AssetsTotal LiabilitiesTotal Liabilities and Stockholders' EquityTotal Long-term InvestmentsTotal Long-term LiabilitiesTotal Paid-in CapitalTotal Paid-in Capital and Retained EarningsTotal Property, Plant and EquipmentTotal Stockholders' Equity

Entry field with correct answer

$

Entry field with incorrect answer

Entry field with correct answer

Entry field with incorrect answer

Entry field with correct answer        Additional Paid-in Capital    Capital Stock    Current Assets    Current Liabilities    Intangible Assets    Long-term Investments    Long-term Liabilities    Property, Plant and Equipment    Stockholders' Equity    Total Assets    Total Capital Stock    Total Current Assets    Total Current Liabilities    Total Intangible Assets    Total Liabilities    Total Liabilities and Stockholders' Equity    Total Long-term Investments    Total Long-term Liabilities    Total Paid-in Capital    Total Paid-in Capital and Retained Earnings    Total Property, Plant and Equipment    Total Stockholders' Equity

Entry field with incorrect answer

Entry field with correct answer Additional Paid-in CapitalCapital StockCurrent AssetsCurrent LiabilitiesIntangible AssetsLong-term InvestmentsLong-term LiabilitiesProperty, Plant and EquipmentStockholders' EquityTotal AssetsTotal Capital StockTotal Current AssetsTotal Current LiabilitiesTotal Intangible AssetsTotal LiabilitiesTotal Liabilities and Stockholders' EquityTotal Long-term InvestmentsTotal Long-term LiabilitiesTotal Paid-in CapitalTotal Paid-in Capital and Retained EarningsTotal Property, Plant and EquipmentTotal Stockholders' Equity

Entry field with correct answer

$

Entry field with incorrect answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with incorrect answer

Entry field with incorrect answer

Entry field with correct answer    Additional Paid-in Capital    Capital Stock    Current Assets    Current Liabilities    Intangible Assets    Long-term Investments    Long-term Liabilities    Property, Plant and Equipment    Stockholders' Equity    Total Assets    Total Capital Stock    Total Current Assets    Total Current Liabilities    Total Intangible Assets    Total Liabilities    Total Liabilities and Stockholders' Equity    Total Long-term Investments    Total Long-term Liabilities    Total Paid-in Capital    Total Paid-in Capital and Retained Earnings    Total Property, Plant and Equipment    Total Stockholders' Equity    

Entry field with incorrect answer

Entry field with correct answer

Entry field with incorrect answer

Entry field with correct answer Additional Paid-in CapitalCapital StockCurrent AssetsCurrent LiabilitiesIntangible AssetsLong-term InvestmentsLong-term LiabilitiesProperty, Plant and EquipmentStockholders' EquityTotal AssetsTotal Capital StockTotal Current AssetsTotal Current LiabilitiesTotal Intangible AssetsTotal LiabilitiesTotal Liabilities and Stockholders' EquityTotal Long-term InvestmentsTotal Long-term LiabilitiesTotal Paid-in CapitalTotal Paid-in Capital and Retained EarningsTotal Property, Plant and EquipmentTotal Stockholders' Equity

Entry field with incorrect answer

Entry field with correct answer AddLess

:

Entry field with correct answer

Entry field with incorrect answer

Entry field with correct answer    Additional Paid-in Capital    Capital Stock    Current Assets    Current Liabilities    Intangible Assets    Long-term Investments    Long-term Liabilities    Property, Plant and Equipment    Stockholders' Equity    Total Assets    Total Capital Stock    Total Current Assets    Total Current Liabilities    Total Intangible Assets    Total Liabilities    Total Liabilities and Stockholders' Equity    Total Long-term Investments    Total Long-term Liabilities    Total Paid-in Capital    Total Paid-in Capital and Retained Earnings    Total Property, Plant and Equipment    Total Stockholders' Equity    

$

Entry field with incorrect answer
Answer #1

NASH COMPANY
Stockholders’ Equity

Entry field with incorrect answer now contains modified data December 31, 2021

Capital stock

Entry field with correct answerPreferred stock

$3548000

Entry field with correct answerCommon stock

Entry field with incorrect answer10185500

Entry field with correct answer Total capital stock

Entry field with incorrect answer13733500

Entry field with correct answer Additional Paid-in Capital

Paid in capital in excess of par- Preferred stockEntry field with correct answer

$294600

Entry field with correct answerPaid in capital in excess of par- Common stock

27215600Entry field with correct answer

Paid in capital from Treasury stockEntry field with correct answer

18400Entry field with incorrect answer

Entry field with incorrect answer27528600

Entry field with correct answer Total Paid-in Capital  

Entry field with incorrect answer41262100

Entry field with correct answerRetained earnings

4200275

Entry field with correct answer Total Paid-in Capital and Retained Earnings

45462375

Less :Treasury stockEntry field with correct answer

-197000Entry field with incorrect answer

Total Stockholders' Equity     $45265375

Preferred stock= $2964000+(29200*$20)= $3548000

Common stock= $9955000+(46100*$5)= $10185500

Paid in capital in excess of par- Preferred stock= $207000+($23-20)*29200= $294600

Paid in capital in excess of par- Common stock= $26478000+($21-5)*46100= $27215600

Paid in capital from Treasury stockEntry field with correct answer= ($12-10)*9200= $18400

Retained earnings= Beginning balance+Net income-Preferred dividend-Common dividend

= $4495000+2138000-283840-2148885= $4200275

Preferred dividend= (148200+29200)*$20*8%= $283840

Common stock shares for calculation of Common stock dividend

Shares= (1991000+46100)*2-28900+9200= 4054500

Common stock dividend= 4054500*0.53= $2148885

Treasury stock= (28900-9200)*$10= $197000